# Your AI, Not Your View: The Bias of LLMs in Investment Analysis

### v1

@@ 3.4. Bias Verification
$\phi_{s}^{\text{int}}=\frac{N_{\text{flip}}^{(s)}}{N}.$
A low $\phi_{s}^{\text{int}}$ implies that the model’s bias overrides qualitatively stronger counter evidence.
These verification methods quantify the transition from mere preference to obdurate bias, highlighting risks in high-stakes applications.
@@ 4. Results
This section presents our empirical findings in sequence with our research questions. First, Section 4.1 addresses RQ1 by identifying the intrinsic preferences of LLMs for stock attributes (sector, size) and investment styles (momentum). To validate these observed differences, we conducted statistical tests to quantify the significance of these preferences. Next, Section 4.2 addresses RQ2 by testing if these preferences become systematic biases under contradictory evidence, using Approach 1 for attributes and Approach 2 for style. Finally, Section 4.3 analyzes the link between preference strength and the model’s internal uncertainty, as measured by entropy.
4.1. Intrinsic Preferences of LLMs
4.1.1. Sector Preference
Our analysis of inherent sector preferences reveals significant variation in both the intensity and range of preferences across the evaluated LLMs (Table 1). For instance, models like Llama4-Scout and DeepSeek-V3 exhibit consistently high preference scores across most sectors, indicating a strong reliance on their internal knowledge representations. In stark contrast, GPT-4.1 and Mistral-24B not only display lower overall preference scores but also show minimal variation between sectors. To quantify these differences, we conducted independent samples t-tests to compare the mean preference scores between high-preference and low-preference sectors for each model (Table LABEL:tab:ttest_sector).
Model High-Pref Low-Pref Diff p-value
Llama4-Scout Energy Consumer Defensive 0.2064 $<0.001^{***}$
DeepSeek-V3 Technology Basic Materials 0.2090 0.014
Qwen3-235B Utilities Consumer Cyclical 0.2361 0.003∗∗
Gemini-2.5 Energy Basic Materials 0.2035 0.035
GPT-4.1 Energy Communication Services 0.1398 0.091
Mistral-24B Basic Materials Communication Services 0.1444 0.124
The t-test results confirm these observations. The preference gaps for Llama4-Scout, Qwen3-235B, DeepSeek-V3, Gemini-2.5 are statistically significant, providing quantitative evidence of a genuine preference. Conversely, the gaps for GPT-4.1 and Mistral-24B are not statistically significant, corroborating that they possess flatter preference landscapes and exhibit a relatively less distinct preference among sectors.
Ultimately, our findings indicate that the strength of preference is a function of the model’s identity rather than any universally preferred sector. This underscores the critical importance of auditing and selecting LLMs for their inherent preferences, especially for deployment in sensitive, real-world applications like financial analysis where model objectivity is paramount.
4.1.2. Size Preference
Model Q1 Q2 Q3 Q4
Llama4-Scout 0.88 0.89 0.82 0.83
DeepSeek-V3 0.86 0.79 0.73 0.67
Qwen3-235B 0.58 0.51 0.49 0.46
Gemini-2.5 0.53 0.46 0.37 0.41
GPT-4.1 0.43 0.43 0.42 0.40
Mistral-24B 0.45 0.44 0.40 0.38
Table 3. Preference scores of each model across four market capitalization quantiles. Q1 corresponds to the highest capitalization quantile and Q4 to the lowest. Overall, the models exhibit a general preference for companies with higher market capitalization.
We investigate whether LLMs exhibit a preference for companies of a certain size, a factor that could influence their outputs in financial applications. To this end, we measure model preference scores across four market capitalization quantiles (Q1: highest, Q4: lowest), with detailed results presented in Table 3. Our findings indicate that most models show a preference for higher-capitalization companies, though the strength of this tendency varies significantly among them. DeepSeek-V3 displays the most pronounced effect, with a strong preference for Q1 that diminishes sharply for lower quantiles. Conversely, GPT-4.1 exhibits nearly uniform preference scores, suggesting its evaluations are largely invariant to company size. Other models, including Gemini-2.5-flash and Qwen3-235B, show a similar but less pronounced downward trend from Q1 to Q4.
To statistically validate these observed trends, we performed an independent samples t-test comparing the preference scores between the highest- and lowest-preference quantiles for each model (Table 4). The analysis confirms that the preference gap is statistically significant for DeepSeek-V3, Llama4-Scout, Qwen3-235B, and Gemini-2.5-flash, revealing a consistent tendency to favor larger companies. In contrast, the preference difference for GPT-4.1 was not statistically significant, corroborating that its judgments are less affected by this factor.
Model High-Pref Low-Pref Diff p-value
Llama4-Scout Q2 Q3 0.0719 0.015
DeepSeek-V3 Q1 Q4 0.1869 $<0.001^{***}$
Qwen3-235B Q1 Q4 0.1178 0.004∗∗
Gemini-2.5 Q1 Q3 0.1514 $<0.001^{***}$
GPT-4.1 Q2 Q4 0.0321 0.417
Mistral-24B Q1 Q4 0.0785 0.054
@@ 4.1.2. Size Preference
Table 4. Independent samples t-test of the preference gap between each model’s highest and lowest preference quantiles. The gap was statistically significant for all models except GPT-4.1 and Mistral-24B.
We attribute this behavior to a popularity effect, wherein greater data volume and richness for larger, well-known corporations in the training corpora lead the models to develop stronger priors for them. This finding has critical implications for the application of LLMs in finance. The models’ inherent inclination towards large-cap stocks could lead to the systematic overlooking of smaller-cap companies, irrespective of their fundamental merits. Therefore, we advise practitioners to be mindful of this characteristic and to account for it when using these models for tasks like portfolio construction.
4.1.3. Momentum Preference
@@
Measuring model preferences for investment styles like momentum or contrarian requires a different approach than the previously discussed sector or size analyses. Unlike a specific sector or size, an investment style can be explicitly framed as $\mathcal{E}_{\text{buy}}^{(s)}$ or $\mathcal{E}_{\text{sell}}^{(s)}$. We leverage this by designing a prompt where $\mathcal{E}_{\text{buy}}^{(s)}$ is based on one view (e.g., momentum), while $\mathcal{E}_{\text{sell}}^{(s)}$ is based on the opposing view (e.g., contrarian). To mitigate potential positional bias, we ensured a balanced experimental design where each view was used to generate both $\mathcal{E}_{\text{buy}}^{(s)}$ and $\mathcal{E}_{\text{sell}}^{(s)}$ an equal number of times.
Model High-Pref Low-Pref Diff p-value
Llama4-Scout contrarian momentum 0.2896 $<0.001^{***}$
DeepSeek-V3 contrarian momentum 0.2541 $<0.001^{***}$
Qwen3-235B contrarian momentum 0.6803 0.037
Gemini-2.5 contrarian momentum 0.0269 0.690
GPT-4.1 contrarian momentum 0.4040 $<0.001^{***}$
Mistral-24B contrarian momentum 0.3056 0.579
Table 5. Chi-Square test of the preference gap between contrarian and momentum views. The gap was statistically significant for all models except for Mistral-24B and Gemini-2.5-flash.
Figure 4. Decision flip rates under varying volumes of evidence for sector and size preferences. The ratios (e.g., 2—3) denote the count of supporting vs. counter- evidences, while the percentages represent the proportion of counter-evidence in the mix. The results reveal a sharp contrast: the flip rate is near 1.0 for all models when only counter-evidence is presented (100% case), but drops significantly the moment any supporting evidence is introduced, indicating a strong difficulty in reversing decisions under conflicting information.
@@
In this setup, if the model ultimately chooses the buy action, the investment view that generated $\mathcal{E}_{\text{buy}}^{(s)}$ is considered to have won. We quantify the model’s preference by repeating this process and calculating the win rate for each investment view.
@@
Figure 3 illustrates the preference of various models for contrarian versus momentum views. Our analysis reveals a consistent preference across all evaluated models toward the contrarian view. In particular, Qwen3-235B exhibits the strongest preference, with a high win rate for a contrarian stance and a correspondingly low rate for momentum. Models such as DeepSeek-V3, Llama4-Scout, and GPT-4.1 also display clear contrarian inclinations, albeit with varying intensities. For Gemini-2.5-flash, the contrarian preference is evident but marginal, with win rates showing a negligible difference between the two views.
@@
To statistically validate these observed tendencies, we performed a Chi-Square test to determine if the difference in win rates between the contrarian and momentum views was significant, with the results presented in Table 5. The analysis confirms that the preference for the contrarian view is statistically significant for DeepSeek-V3, Llama4-Scout, Qwen3-235B, and GPT-4.1. In contrast, for Mistral-24B and Gemini-2.5, the difference in preference is not statistically significant, suggesting their observed tendencies may be due to chance.
These results highlight the need to account for intrinsic view preferences in LLMs, as they may cause discrepancies between anticipated and actual outputs in decision-making tasks.
4.2. From Preference to Bias: An Experimental Validation
@@
This section examines the Decision Flip Rate, a metric quantifying the resilience of a model’s initial preference when exposed to a high proportion of counter-evidence. The experiments aim to assess the extent of confirmation bias, with reported values indicating the frequency of decision reversals under controlled conditions where evidence is intentionally skewed against the model’s preference.
Approach 1: Verification by Evidence Volume
@@
This experiment was designed to observe how much an initial decision is reversed when a model is provided with a weighted amount of evidence that opposes its existing preference. This rate of change, measured as $\phi_{s}^{\text{vol}}$, probes the persistence of bias, yielding results consistent with prior observations of contradictory LLM behaviors (Xie et al., 2023). Figure 4 presents the $\phi_{s}^{\text{vol}}$ values across models for different evidence ratios, where a ratio (e.g., 2—3) denotes the volume of supporting and counter-evidence, respectively.
@@
When provided only with counter-evidence, all models exhibited high receptivity, overriding their internal knowledge and achieving $\phi_{s}^{\text{vol}}$ values near 1.0. However, in situations where supporting and counter-evidence were mixed, creating a knowledge conflict, the $\phi_{s}^{\text{vol}}$ values dropped sharply. This phenomenon occurred despite the amount of counter-evidence always being greater than the supporting evidence in all experimental conditions, strongly suggesting that models selectively adhere to information that aligns with their pre-existing inclinations.
@@
This rigidity was more evident in models with strong inherent preferences. For instance, Llama4-Scout and DeepSeek-V3, which had high preference scores across sectors, recorded particularly low $\phi_{s}^{\text{vol}}$ values. These models struggled to reverse their decisions, especially when the volume difference between supporting and counter-evidence was small. Similarly, Qwen3-235b also showed reduced flexibility at lower proportions of counter-evidence.
@@
In contrast, models with overall lower preference scores demonstrated greater adaptability. GPT-4.1 and Gemini-2.5-flash maintained higher $\phi_{s}^{\text{vol}}$ values, remaining relatively responsive even when the difference in evidence volume was minimal. Although their $\phi_{s}^{\text{vol}}$ values fell short of expectations despite the counter-evidence majority, this pattern shows a direct correlation with initial preference strength. In other words, the stronger a model’s inherent bias, the more its stubbornness is amplified when the difference in the volume of supporting and counter-evidence is small. Consequently, this finding suggests a significant risk in real-world financial contexts where conflicting information is present (for instance, when price indicators are negative but related news is positive). In such cases, a model could trust only one side of the evidence due to its inherent bias, leading to flawed judgments.
Approach 2: Verification by Evidence Intensity
@@
This approach investigates model sensitivity by maintaining volumetric parity while escalating the intensity increment, $\Delta$, of the counter-evidence. Figure 5 plots the intensity-driven flip rate ($\phi_{s}^{\text{int}}$) against $\Delta$ values of 1, 3, 5, and 10. The results delineate a clear sensitivity spectrum among the models, which correlates with the prior View Preferences analysis.
@@
While the graph shows a gradual upward trend in $\phi_{s}^{\text{int}}$ for all models as $\Delta$ increases, the more notable finding lies in the magnitude of this increase and the final values. Even when presented with very strong counter-evidence ($\Delta=10$), the majority of models recorded low $\phi_{s}^{\text{int}}$ values below 60%. This signifies that the models’ confirmation bias is not easily overcome, even by qualitatively superior counter-evidence.
Amidst this overall rigidity, a distinct performance gap emerged based on the models’ initial preference strengths. Exhibiting the most balanced preference profile, Gemini-2.5-flash recorded the highest $\phi_{s}^{\text{int}}$ and showed a stark contrast to all other models. This clearly demonstrates that an absence of strong initial bias leads to greater flexibility.
@@
Conversely, models identified with stronger and more polarized initial preferences formed the lower-performing group. These models consistently recorded low $\phi_{s}^{\text{int}}$ values, signifying a more stubborn confirmation bias. The behavior of Qwen3-235B, which had one of the largest preference gaps, exemplifies this resistance, as it remains one of the least likely models to reverse its decision even when the counter-evidence is significantly more intense.
@@
Synthesizing these results provides a deeper insight into model behavior. Even when presented with qualitatively superior counter-evidence ($\Delta=10$), models show a strong tendency to struggle with decision reversal due to their initial preferences. This rigidity poses a tangible risk when considering the findings from our prior analysis, where all models commonly preferred a contrarian view over a momentum view. It implies that a model’s inherent bias toward a specific investment perspective could cause it to ignore or undervalue strong opposing evidence, potentially leading to skewed conclusions.
@@
Figure 5. Decision flip rates under varying volumes of evidence for momentum preferences. Even as counter-evidence intensity ($\Delta$) increases, the decision flip rate ($\phi_{s}^{\text{int}}$) for most models remains low, indicating strong confirmation bias. Gemini-2.5-flash, which had the least initial bias, shows the most flexible response, demonstrating that initial preference is a key predictor of bias.

### v4

@@ 3.4.2. Approach 2: Verification by Evidence Intensity
(5) $\phi_{s^{*}}^{\text{int}}=\frac{N_{\text{flip}}^{(s^{*})}}{N}.$
A low $\phi_{s^{*}}^{\text{int}}$ implies that the model’s intrinsic bias is sufficiently tenacious to override even qualitatively stronger counter-evidence.
@@ 4. Results
This section presents our empirical findings, structured to answer our research questions sequentially. First, to address RQ1, Section 4.1 identifies and quantifies the LLM’s latent biases toward factors like sector, size, and momentum, confirming their statistical significance. Next, Section 4.2 addresses RQ2 by quantitatively analyzing the resulting bias. We demonstrate how these biases lead to a stubborn refusal to revise judgments when the model is challenged with stronger counter-evidence in terms of both volume and intensity. Finally, Section 4.3 further investigates this phenomenon by linking the strength of the bias to the model’s internal uncertainty, as measured by entropy.
4.1. Intrinsic Bias of LLMs
4.1.1. Sector Bias
The bias scores presented in this section were calculated over 10 trials for each stock. To ensure the robustness of our findings against the inherent variability of LLM responses, we conducted this entire process in three independent sets; all results are reported as a mean score with the corresponding standard deviation. Our analysis of inherent sector bias reveals two notable patterns: a prevalent bias toward the Technology sector and an overall tendency for positive bias scores, reflecting a default inclination to buy. However, the strength of this bias varies considerably across models (Figure 3). For example, models such as Llama4-Scout and DeepSeek-V3 follow the general trend, consistently exhibiting high positive bias scores across most sectors. In contrast, other models display more nuanced behavior; GPT-4.1, for instance, not only shows a lower average bias but also exhibits negative bias scores in certain sectors, deviating from the common buy-oriented tendency.
To quantify these differences, we conducted independent samples t-tests comparing the mean bias scores between each model’s highest and lowest bias sectors (Table 1). Crucially, the analysis confirms that the gap between the most and least favored sectors is statistically significant for every model evaluated. This universal finding provides strong evidence that all models possess distinct and deeply embedded sector biases in their knowledge representations.
@@ 4.1.1. Sector Bias
Table 1. Independent-samples t-test results for the score gap between each model’s highest- and lowest-scoring sectors.
Model High-Score Low-Score p-value
Llama4-Scout Energy Consumer Defensive $0.001^{***}$
DeepSeek-V3 Technology Financial Services $0.001^{***}$
Gemini-2.5-flash Technology Financial Services $0.001^{***}$
Qwen3-235B Technology Consumer Defensive $0.001^{***}$
Mistral-Small Technology Consumer Defensive $0.001^{***}$
GPT-4.1 Technology Consumer Defensive $0.001^{***}$
$^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$
4.1.2. Size Bias
We investigate whether LLMs exhibit a bias for companies of a certain size, a factor that could influence their outputs in financial applications. To this end, we measure model bias scores across four market capitalization quantiles (Q1: highest, Q4: lowest), with detailed results presented in Figure 4. Across all models examined, we observe a consistent pattern: bias scores generally decline as company size decreases, indicating a stronger preference for large-capitalization firms that weakens toward smaller ones. Notably, one model’s scores even become negative in the lowest two quantiles (Q3 and Q4), suggesting not merely a loss of preference but at times even a negative stance toward smaller-capitalization companies. These results highlight that size bias is a pervasive feature among LLMs, and its intensity and direction can differ substantially depending on the model.
Figure 4. Size bias scores for each evaluated LLM across four market-capitalization quantiles (Q1: largest, Q4: smallest). Bias scores consistently decline as company size decreases.
To statistically validate these observed trends, we performed independent-samples t-tests comparing the bias scores between the highest- and lowest-bias quantiles for each model (Table 2). The results show that the gap in bias scores is statistically significant across all models evaluated, confirming that size-related bias is a consistent characteristic of LLMs. Although the magnitude varies, every model shows a clear size-related difference in bias scores.
Table 2. Independent-samples t-test results for the score gap between each model’s highest- and lowest-scoring quantiles.
Model High-Score Low-Score p-value
Llama4-Scout Q1 Q3 $0.001^{***}$
DeepSeek-V3 Q1 (Largest) Q4 (Smallest) $0.001^{***}$
Gemini-2.5-flash Q1 Q4 $0.001^{***}$
Qwen3-235B Q1 Q4 $0.001^{***}$
Mistral-Small Q1 Q3 $0.001^{***}$
GPT-4.1 Q1 Q4 $0.001^{***}$
We attribute these biases to underlying popularity effects, wherein sectors and companies with greater market prominence—such as well-known industries or large-cap firms—are likely to appear more frequently and in richer contexts within the training corpora. As a result, LLMs tend to develop stronger priors for these dominant sectors and larger companies. This dual bias toward popular sectors and high-capitalization stocks has critical implications for financial applications: it can lead models to systematically overvalue certain industries and large-cap firms while undervaluing less represented sectors or smaller-cap companies, regardless of their fundamental merits. Practitioners should remain mindful of these tendencies and implement appropriate auditing or corrective measures when deploying LLMs for tasks such as portfolio construction.
4.1.3. Momentum Bias
@@
Measuring model bias for investment styles like momentum or contrarian requires a different approach than the previously discussed sector or size analyses. Unlike a specific sector or size, an investment style can be explicitly framed as $\mathcal{E}_{\text{buy}}^{(s)}$ or $\mathcal{E}_{\text{sell}}^{(s)}$. We leverage this by designing a prompt where $\mathcal{E}_{\text{buy}}^{(s)}$ is based on one view (e.g., momentum), while $\mathcal{E}_{\text{sell}}^{(s)}$ is based on the opposing view (e.g., contrarian). To mitigate potential positional bias, we ensured a balanced experimental design where each view was used to generate both $\mathcal{E}_{\text{buy}}^{(s)}$ and $\mathcal{E}_{\text{sell}}^{(s)}$ an equal number of times.
@@
In this setup, if the model ultimately chooses the buy action, the investment view that generated $\mathcal{E}_{\text{buy}}^{(s)}$ is considered to have won. We quantify the model’s bias by repeating this process and calculating the win rate for each investment view.
Figure 6. Decision flip rates under varying volumes of evidence for sector and size bias. The ratios (e.g., 2|3) denote the count of supporting vs. counter- evidence, while the percentages represent the proportion of counter-evidence in the mix. The results reveal a sharp contrast: the flip rate is near 1.0 for all models when only counter-evidence is presented (100% case), but drops significantly the moment any supporting evidence is introduced, indicating a strong difficulty in reversing decisions under conflicting information.
@@
Figure 5 illustrates the bias of various models for contrarian versus momentum views. Our analysis reveals a consistent preference across all evaluated models toward the contrarian view. In particular, Qwen3-235B exhibits the strongest preference, with a high win rate for a contrarian stance and a correspondingly low rate for momentum. Models such as DeepSeek-V3, Llama4-Scout, and GPT-4.1 also display clear contrarian inclinations, albeit with varying intensities. For Gemini-2.5-flash, the contrarian preference is evident but marginal, with win rates showing a negligible difference between the two views.
@@
To statistically validate these observed tendencies, we performed a Chi-Square test to assess whether the difference in win rates between the contrarian and momentum views was significant (Table 3). The results show that all models exhibit a statistically significant preference for the contrarian view, confirming that this tendency is consistent across the evaluated LLMs.
Table 3. Chi-Square test results for the score gap between contrarian and momentum views.
Model High-Score Low-Score p-value
Llama4-Scout contrarian momentum $0.001^{***}$
DeepSeek-V3 contrarian momentum $0.001^{***}$
Gemini-2.5-flash contrarian momentum $0.0052^{**}$
Qwen3-235B contrarian momentum $0.001^{***}$
Mistral-Small contrarian momentum $0.001^{***}$
GPT-4.1 contrarian momentum $0.001^{***}$
These results highlight the importance of accounting for intrinsic view preferences in LLMs. A consistent bias toward the contrarian style can skew investment decisions, favoring underperforming assets even when market signals support momentum strategies. Recognizing these biases is essential for deploying LLMs in finance, where unmitigated view preferences may lead to unintended portfolio imbalances. Practitioners should audit models for these tendencies and consider mitigation strategies to ensure decisions align with intended investment objectives.
4.2. Bias Verification with Counter-Evidence
@@
This section examines the Decision Flip Rate, a metric quantifying the resilience of a model’s initial bias when exposed to a high proportion of counter-evidence. The experiments aim to assess the resilience of this bias, with reported values indicating the frequency of decision reversals under controlled conditions where evidence is intentionally skewed against the model’s established bias.
4.2.1. Approach 1: Verification by Evidence Volume
@@
This experiment was designed to observe how much an initial decision is reversed when a model is provided with a weighted amount of evidence that opposes its existing bias. This rate of change, measured as $\phi_{s^{*}}^{\text{vol}}$, probes the persistence of bias, yielding results consistent with prior observations of contradictory LLM behaviors (Xie et al., 2023). Figure 6 presents the $\phi_{s^{*}}^{\text{vol}}$ values for each model at various ratios of supporting versus counter-evidence (e.g., 2|3).
@@
When provided only with counter-evidence, all models exhibited high receptivity, overriding their internal knowledge and achieving $\phi_{s^{*}}^{\text{vol}}$ values near 1.0. However, in situations where supporting and counter-evidence were mixed, creating a knowledge conflict, the $\phi_{s^{*}}^{\text{vol}}$ values dropped sharply. This phenomenon occurred despite the amount of counter-evidence always being greater than the supporting evidence in all experimental conditions, strongly suggesting that models selectively adhere to information that aligns with their pre-existing inclinations.
@@
This rigidity was more evident in models with strong inherent bias. For instance, Llama4-Scout and DeepSeek-V3, which had high bias scores across sectors, recorded particularly low $\phi_{s^{*}}^{\text{vol}}$ values. These models struggled to reverse their decisions, especially when the volume difference between supporting and counter-evidence was small. Similarly, Qwen3-235b also showed reduced flexibility at lower proportions of counter-evidence.
@@
In contrast, models with overall lower bias scores demonstrated greater adaptability. GPT-4.1 and Gemini-2.5-flash maintained higher $\phi_{s}^{\text{vol}}$ values, remaining relatively responsive even when the difference in evidence volume was minimal. Although their $\phi_{s^{*}}^{\text{vol}}$ values fell short of expectations despite the counter-evidence majority, this pattern shows a direct correlation with initial bias strength. In other words, the stronger a model’s inherent bias, the more its stubbornness is amplified when the difference in the volume of supporting and counter-evidence is small. Consequently, this finding suggests a significant risk in real-world financial contexts where conflicting information is present (for instance, when price indicators are negative but related news is positive). In such cases, a model could trust only one side of the evidence due to its inherent bias, leading to flawed judgments.
4.2.2. Approach 2: Verification by Evidence Intensity
@@
This approach investigates model sensitivity by maintaining volumetric parity while escalating the intensity increment, $\Delta$, of the counter-evidence. Figure 7 plots the intensity-driven flip rate ($\phi_{s^{*}}^{\text{int}}$) against $\Delta$ values of 1, 3, 5, and 10. The results delineate a clear sensitivity spectrum among the models, which correlates with the prior view bias analysis.
@@
While the graph shows a gradual upward trend in $\phi_{s^{*}}^{\text{int}}$ for all models as $\Delta$ increases, the more notable finding lies in the magnitude of this increase and the final values. Even when presented with very strong counter-evidence ($\Delta=10$), the majority of models recorded low $\phi_{s^{*}}^{\text{int}}$ values below 60%. This signifies that the models’ confirmation bias is not easily overcome, even by qualitatively superior counter-evidence.
Finally, we observe a strong correlation between bias strength and model rigidity. Notably, Gemini-2.5-flash, the model with the most balanced bias profile, achieves a substantially higher flip rate ($\phi_{s^{*}}^{\text{int}}$) than its counterparts. This finding suggests that a low initial bias is a key predictor of a model’s ability to adapt when presented with qualitatively superior counter-evidence.
@@
Conversely, models identified with stronger and more polarized bias formed the lower-performing group. These models consistently recorded low $\phi_{s^{*}}^{\text{int}}$ values, signifying a more stubborn confirmation bias. The behavior of Qwen3-235B, which had one of the largest bias gaps, exemplifies this resistance, as it remains one of the least likely models to reverse its decision even when the counter-evidence is significantly more intense.
@@
Synthesizing these results provides a deeper insight into model behavior. Even when presented with qualitatively superior counter-evidence ($\Delta=10$), models show a strong tendency to struggle with decision reversal due to their bias. This rigidity poses a tangible risk when considering the findings from our prior analysis, where all models commonly biased a contrarian view over a momentum view. It implies that a model’s inherent bias toward a specific investment perspective could cause it to ignore or undervalue strong opposing evidence, potentially leading to skewed conclusions.
@@
Figure 7. Decision flip rates under varying intensities of evidence for momentum bias. Even as counter-evidence intensity ($\Delta$) increases, the decision flip rate ($\phi_{s}^{\text{int}}$) for most models remains low, indicating strong confirmation bias. Gemini-2.5-flash, which had the least initial bias, shows the most flexible response.